OpenStudy (anonymous):

Write an equation in point-slope form for the point (3, -4) and slope of .

OpenStudy (lgbasallote):

slope of?

OpenStudy (anonymous):

1/3

OpenStudy (lgbasallote):

point slope form: \(\LARGE y - y_{1} = m(x - x_{1})\)

OpenStudy (anonymous):

so what would i do from there?

OpenStudy (lgbasallote):

substitute then you're all set..though you might wanna transform it to the form Ax + By + C = 0

OpenStudy (anonymous):

how do i sub it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y - y1 = m(x - x1) x1 = 3, y1 = -4, m = ? You substitute all these numbers into the formula y - (-4) = m(x - 3) y + 4 = m(x - 3) whatever your slope is, then sub it in for the m and solve.

OpenStudy (anonymous):

so what would my final answer be